. Truck #2 has a list price of $16,000 and is acquired for a down payment of $2,000 cash and a zero- interest-bearing note with a face amount of $14,000. The note is due April 1, 2015. Clarkson would normally have to pay interest at a rate of 10% for such a borrowing, and the dealership has an incremental borrowing rate of 8%. Prepare the appropriate journal entries for the above transactions for Clarkson Corporation Please explain answer
Answer #1
Calculating present value of principal amount to be paid after 1 year
Present value= Principal value x PVF @ 10% for 1 year
Present value= 14000 x .909= 12726
Discount on note payable = 14000 - 12726= 1274
Value of truck shall be =2000 +12726 =14726
Journal entry
Debit Credit
Truck 14726
Discount on notes payable 1274
cash 2000
notes payable 14000
( Being asset recorded)